# Extending Dense Passage Retrieval with Temporal Information

### v1

@@ 4.5. Settings
The TempDPR model was trained using an in-batch negative sampling approach with a batch size of 32. For our experiments, we trained the TempDPR for $5$ epochs with a learning rate set to $1\times 10^{-5}$, using the Adam optimizer and a warm-up ratio of $0.1$. All experiments (except Section 5.4) were conducted with 4 randomly selected negative passages per query, ensuring that none of the negative passages contained the correct answer.
@@ 5.1. Impact of Temporal Fusion Strategies
We first start with investigating the performance of different temporal fusion techniques in TempDPR model, such as Feature Stacking (FS), Vector Summation (VS), Relative Embeddings (RE), and Element-Wise Interaction (EWI) with state-of-the-art temporal retrieval models, BiTimeBert (Wang et al., 2023) and TS-Retriever (Wu et al., 2024), as well as the baseline VanillaDPR (Karpukhin et al., 2020). The results in Table 2 highlight the effectiveness of our temporal fusion techniques across both the ChroniclingAmericaQA and ArchivalQA datasets. On the ChroniclingAmericaQA dataset, TempDPR FS outperforms all other models, achieving a Top-1 accuracy of 50.22% on the validation set. This represents a substantial improvement over BiTimeBert, which achieves 35.74%, and TS-Retriever, which reaches 47.85%. In terms of Top-100 accuracy, TempDPR FS achieves 86.22%, surpassing TS-Retriever and BiTimeBert by 2.79 and 7.91%, respectively. Similarly, on the ArchivalQA dataset, TempDPR FS achieves state-of-the-art results, with a Top-1 accuracy of 69.72% on the validation set. This is a clear improvement over BiTimeBert (61.71%) and TS-Retriever (63.44%). For higher $k$ values, such as the Top-100 accuracy, TempDPR FS reaches 97.77%, representing a notable improvement over both TS-Retriever (97.33%) and BiTimeBert (95.52%). These results demonstrate the capability of TempDPR to capture and align temporal signals more effectively than competing methods.
@@ 5.2. Retrieval Performance Analysis
In this section, we present a comprehensive comparison of TempDPR with baseline methods, focusing on its overall retrieval performance across various metrics, such as top-$k$ accuracy, nDCG, and MAP. We evaluate TempDPR alongside VanillaDPR, DateAsTag, and DateAsToken models.
Table 3 shows the retrieval performance of TempDPR against three baselines on the ArchivalQA and ChroniclingAmericaQA datasets, using top-$k$ accuracy ($k\in\{1,5,10,20,50,100\}$). Across both datasets, TempDPR outperforms the VanillaDPR, DateAsTag, and DateAsToken models, demonstrating the benefit of directly incorporating temporal embeddings into the retrieval process. On ChroniclingAmericaQA, TempDPR achieves a top-100 accuracy of 86.22% on the validation set, compared to 83.53% for VanillaDPR. Table 4 summarises the nDCG and MAP metrics. TempDPR achieves an nDCG@10 of 53.03 on the validation set of ChroniclingAmericaQA, outperforming the DateAsTag and DateAsToken models by over $4\%$. MAP shows a similar trend, with TempDPR achieving a MAP@20 of 53.88% on the test set, significantly higher than the 49.74% obtained by the VanillaDPR baseline.
Similar trends are observed on ArchivalQA, where TempDPR demonstrates superior performance in ranking. As shown in Table 4, TempDPR achieves nDCG@10 score of 67.76 on the validation set, surpassing the DateAsTag and DateAsToken approaches, which reach 54.71 and 60.71, respectively. Also, TempDPR outperforms the other methods significantly, achieving a MAP@10 of 63.64 on the validation set. The DateAsTag and DateAsToken models achieve MAP of 50.38 and 56.30, respectively. These results highlight that TempDPR not only retrieves semantically relevant passages but also ranks them more accurately according to their temporal relevance, resulting in improvement compared to baseline methods.
To validate TempDPR ’s performance improvements, we conducted paired t-tests comparing Top-1, Top-50, and Top-100 accuracies with baselines (VanillaDPR, DateAsTag, DateAsToken). Table 6 shows the p-values for these comparisons, all significantly below 0.05, confirming the statistical significance of TempDPR ’s results. For instance, the p-values for Top-1 accuracy were $1.39\times 10^{-5}$, $4.18\times 10^{-7}$, and $2.13\times 10^{-5}$ for VanillaDPR, DateAsTag, and DateAsToken, respectively. Similar trends were observed for Top-50 and Top-100 metrics, demonstrating the robustness of TempDPR ’s improvements across retrieval metrics.
Table 6. Paired t-test results for Top-$k$ accuracy comparisons between TempDPR and baselines (VanillaDPR, DateAsTag, and DateAsToken).
@@ 5.3. Enhancing Temporal Models with Fusion Strategies
To demonstrate the flexibility and impact of TempDPR ’s fusion techniques, we evaluate their ability to enhance the performance of existing temporal retrieval models, including BiTimeBERT (Wang et al., 2023) and TS-Retriever (Wu et al., 2024). By integrating temporal embeddings and leveraging fusion strategies, we showcase how TempDPR can amplify the strengths of SOTA temporal models.
@@
Similar trends are observed on the ArchivalQA dataset. BiTimeBERT alone achieves a Top-1 accuracy of 61.71%, but this increases to 64.93% with TempDPR +${}_{FS}+\texttt{BiTimeBERT}$. TS-Retriever shows an even more pronounced improvement, with its Top-1 accuracy rising from 63.44% to 69.94% when combined with TempDPR and feature stacking.
@@ 5.4. Negative Document Selection Strategies
Table 7. Evaluation of Different Modes on ChroniclingAmericaQA Dataset. The table shows Top-k, nDCG, and MAP metrics for both Validation and Test sets.
@@ 6. Implicit Temporal Questions
TempDPR VS 45.00 67.30 74.70 80.70 86.40 90.10 41.70 66.40 73.80 80.00 85.10 89.60
TempDPR EWI 43.50 65.60 72.40 77.70 84.80 88.50 40.20 63.70 70.40 76.60 83.80 88.00
TempDPR FS 44.80 67.40 74.10 79.60 86.50 90.20 43.30 66.40 74.50 79.60 86.10 89.70
@@ 6.1. Implicit Temporal Questions in Retrieval
The evaluation on the ArchivalQA dataset demonstrates the effectiveness of the proposed TempDPR framework in addressing implicit temporal questions, as shown in Table 8 and Figure 3. TempDPR consistently outperforms baseline models and pretrained temporal models across validation and test sets. For example, VanillaDPR achieves a Top-1 accuracy of 43.10% and a Top-100 accuracy of 89.50% on the validation set, while DateAsToken slightly improves with a Top-100 accuracy of 89.00%. Among temporal models, BiTimeBERT performs the worst (Top-1: 27.70%, Top-100: 81.80%), highlighting its limitations in reasoning about implicit temporal information. TS-Retriever performs better, reaching a Top-1 accuracy of 39.60% and a Top-100 accuracy of 90.90%.
TempDPR models, particularly TempDPR FS, achieve state-of-the-art results, with a Top-1 accuracy of 44.80% and a Top-100 accuracy of 90.20%. TempDPR VS achieves the highest Top-1 accuracy of 45.00%. Ranking quality metrics (Figure 3) further demonstrate TempDPR ’s ability to prioritize relevant passages. For instance, at nDCG@1, VanillaDPR achieves 28.30%, which improves to 29.60% and 30.30% with TempDPR FS and TempDPR VS, respectively. This trend persists across all $k$-values, with TempDPR FS achieving the highest scores, such as 45.09% at nDCG@100, compared to 43.78% for VanillaDPR and 29.29% for BiTimeBERT.
@@ 6.2. Addressing Different Question Types
We next explain how to construct a unified framework for handling different question types. Temporal retrieval systems must address diverse question types, including explicit temporal, implicit temporal, and non-temporal queries. As illustrated in Fig. 4, a query routing mechanism can be implemented to dynamically classify and assign questions to the appropriate retriever. Explicit temporal questions, which contain direct references to dates, are routed to TempDPR, leveraging its ability to integrate temporal embeddings and fusion techniques.
Implicit temporal questions that lack explicit date mentions are routed to the query date prediction model, which predicts a temporal context based on the query’s content. The predicted date is then used to enhance the retrieval process through TempDPR, aligning the inferred temporal context with the document timestamps. Finally, non-temporal queries, which do not involve temporal elements are directed to a general retriever like VanillaDPR. This simple routing framework optimizes retrieval performance across diverse query types by aligning each query with the most effective retriever.
@@ 7. RAG for Temporal Question Answering
The integration of TempDPR into the Retrieval-Augmented Generation (RAG) system involves replacing the traditional retriever component with TempDPR, which retrieves the top-k passages most relevant to a given query by considering both semantic and temporal relevance. These passages are then passed to a pre-trained large language model (LLM), such as Llama-3-8B (AI@Meta, 2024), for answer generation. The pipeline includes three main steps: encoding the query with its temporal context to create dense representations, ranking passages based on their combined semantic and temporal alignment, and generating answers using the retrieved passages as input for the LLM.
For the experiments, Llama-3-8B was used as the LLM, and its performance was evaluated on the ChroniclingAmericaQA and ArchivalQA datasets. The retriever retrieved top-1 passages for each query, which were fed into the LLM for supporting answer generation. Metrics such as Exact Match (EM), Recall (R), and Contains (Con) were used to compare the performance of RAG systems utilizing TempDPR, VanillaDPR, and other baseline models.
Table 9 presents the results of the RAG framework for temporal question answering on explicit datasets. TempDPR consistently outperformed other retrieval methods across all evaluated models and datasets. For example, when paired with Llama-3-8B, TempDPR achieved an EM score of 16.90% and a recall of 28.52% on the validation set of ChroniclingAmericaQA. These scores were significantly higher than those obtained by the DateAsTag and DateAsToken models, which, while showing improvements over VanillaDPR, fell short of TempDPR ’s performance. The DateAsToken model achieved an EM score of 16.57% on the same dataset, slightly below TempDPR.
Across different LLMs, TempDPR provided superior inputs for the generation task, as reflected in its higher metrics. For instance, with Llama-3-8B, TempDPR achieved an EM score of 19.65% on ChroniclingAmericaQA and 45.21% on ArchivalQA, demonstrating its robustness in handling temporally-sensitive queries. These results underline TempDPR’s ability to deliver temporally-aligned and contextually-relevant passages, enabling more accurate and effective answer generation in RAG systems.
@@ 8. Conclusion
In this paper, we introduced TempDPR, a novel framework for Temporal Information Retrieval that integrates temporal information into dense passage retrieval using fusion techniques. By incorporating temporal signals directly into query and document representations, TempDPR bridges the gap between semantic and temporal alignment, achieving superior performance across multiple retrieval metrics. Experiments on ArchivalQA and ChroniclingAmericaQA datasets demonstrate that TempDPR significantly outperforms state-of-the-art baselines, including VanillaDPR, DateAsToken, and DateAsTag models, as well as temporal retrieval models like BiTimeBERT and TS-Retriever.
We also evaluated TempDPR in a Retrieval-Augmented Generation (RAG) setup, where it provided high-quality inputs to large language models, leading to improved answer generation in temporal question answering. Looking forward, we plan to enhance TempDPR’s handling of implicit temporal queries by integrating more advanced query date prediction models.

### v2

@@ 4.5. Settings
The TempRetriever model was trained using an in-batch negative sampling approach with a batch size of 32. For our experiments, we trained the TempRetriever for $5$ epochs with a learning rate set to $1\times 10^{-5}$, using the Adam optimizer and a warm-up ratio of $0.1$. All experiments (except Section 5.4) were conducted with 4 randomly selected negative passages per query, ensuring that none of the negative passages contained the correct answer.
@@ 5.1. Impact of Temporal Fusion Strategies
We first start with investigating the performance of different temporal fusion techniques in TempRetriever model, such as Feature Stacking (FS), Vector Summation (VS), Relative Embeddings (RE), and Element-Wise Interaction (EWI) with state-of-the-art temporal retrieval models, BiTimeBert (Wang et al., 2023) and TS-Retriever (Wu et al., 2024), as well as the baseline VanillaDPR (Karpukhin et al., 2020). The results in Table 2 highlight the effectiveness of our temporal fusion techniques across both the ChroniclingAmericaQA and ArchivalQA datasets. On the ChroniclingAmericaQA dataset, TempRetriever FS outperforms all other models, achieving a Top-1 accuracy of 50.22% on the validation set. This represents a substantial improvement over BiTimeBert, which achieves 35.74%, and TS-Retriever, which reaches 47.85%. In terms of Top-100 accuracy, TempRetriever FS achieves 86.22%, surpassing TS-Retriever and BiTimeBert by 2.79 and 7.91%, respectively. Similarly, on the ArchivalQA dataset, TempRetriever FS achieves state-of-the-art results, with a Top-1 accuracy of 69.72% on the validation set. This is a clear improvement over BiTimeBert (61.71%) and TS-Retriever (63.44%). For higher $k$ values, such as the Top-100 accuracy, TempRetriever FS reaches 97.77%, representing a notable improvement over both TS-Retriever (97.33%) and BiTimeBert (95.52%). These results demonstrate the capability of TempRetriever to capture and align temporal signals more effectively than competing methods.
@@ 5.2. Retrieval Performance Analysis
In this section, we present a comprehensive comparison of TempRetriever with baseline methods, focusing on its overall retrieval performance across various metrics, such as top-$k$ accuracy, nDCG, and MAP. We evaluate TempRetriever alongside VanillaDPR, DateAsTag, and DateAsToken models.
Table 3 shows the retrieval performance of TempRetriever against three baselines on the ArchivalQA and ChroniclingAmericaQA datasets, using top-$k$ accuracy ($k\in\{1,5,10,20,50,100\}$). Across both datasets, TempRetriever outperforms the VanillaDPR, DateAsTag, and DateAsToken models, demonstrating the benefit of directly incorporating temporal embeddings into the retrieval process. On ChroniclingAmericaQA, TempRetriever achieves a top-100 accuracy of 86.22% on the validation set, compared to 83.53% for VanillaDPR. Table 4 summarises the nDCG and MAP metrics. TempRetriever achieves an nDCG@10 of 53.03 on the validation set of ChroniclingAmericaQA, outperforming the DateAsTag and DateAsToken models by over $4\%$. MAP shows a similar trend, with TempRetriever achieving a MAP@20 of 53.88% on the test set, significantly higher than the 49.74% obtained by the VanillaDPR baseline.
Similar trends are observed on ArchivalQA, where TempRetriever demonstrates superior performance in ranking. As shown in Table 4, TempRetriever achieves nDCG@10 score of 67.76 on the validation set, surpassing the DateAsTag and DateAsToken approaches, which reach 54.71 and 60.71, respectively. Also, TempRetriever outperforms the other methods significantly, achieving a MAP@10 of 63.64 on the validation set. The DateAsTag and DateAsToken models achieve MAP of 50.38 and 56.30, respectively. These results highlight that TempRetriever not only retrieves semantically relevant passages but also ranks them more accurately according to their temporal relevance, resulting in improvement compared to baseline methods.
To validate TempRetriever ’s performance improvements, we conducted paired t-tests comparing Top-1, Top-50, and Top-100 accuracies with baselines (VanillaDPR, DateAsTag, DateAsToken). Table 6 shows the p-values for these comparisons, all significantly below 0.05, confirming the statistical significance of TempRetriever ’s results. For instance, the p-values for Top-1 accuracy were $1.39\times 10^{-5}$, $4.18\times 10^{-7}$, and $2.13\times 10^{-5}$ for VanillaDPR, DateAsTag, and DateAsToken, respectively. Similar trends were observed for Top-50 and Top-100 metrics, demonstrating the robustness of TempRetriever ’s improvements across retrieval metrics.
Table 6. Paired t-test results for Top-$k$ accuracy comparisons between TempRetriever and baselines (VanillaDPR, DateAsTag, and DateAsToken).
@@ 5.3. Enhancing Temporal Models with Fusion Strategies
To demonstrate the flexibility and impact of TempRetriever ’s fusion techniques, we evaluate their ability to enhance the performance of existing temporal retrieval models, including BiTimeBERT (Wang et al., 2023) and TS-Retriever (Wu et al., 2024). By integrating temporal embeddings and leveraging fusion strategies, we showcase how TempRetriever can amplify the strengths of SOTA temporal models.
@@
Similar trends are observed on the ArchivalQA dataset. BiTimeBERT alone achieves a Top-1 accuracy of 61.71%, but this increases to 64.93% with TempRetriever +${}_{FS}+\texttt{BiTimeBERT}$. TS-Retriever shows an even more pronounced improvement, with its Top-1 accuracy rising from 63.44% to 69.94% when combined with TempRetriever and feature stacking.
@@ 5.4. Negative Document Selection Strategies
Table 7. Evaluation of Different Negative sampling Modes on ArchivalQA and ChroniclingAmericaQA Datasets. The table shows Top-k, nDCG, and MAP metrics for both Validation and Test sets.
@@ 6. Implicit Temporal Questions
TempRetriever VS 45.00 67.30 74.70 80.70 86.40 90.10 41.70 66.40 73.80 80.00 85.10 89.60
TempRetriever EWI 43.50 65.60 72.40 77.70 84.80 88.50 40.20 63.70 70.40 76.60 83.80 88.00
TempRetriever FS 44.80 67.40 74.10 79.60 86.50 90.20 43.30 66.40 74.50 79.60 86.10 89.70
@@ 6.1. Implicit Temporal Questions in Retrieval
The evaluation on the ArchivalQA dataset demonstrates the effectiveness of the proposed TempRetriever framework in addressing implicit temporal questions, as shown in Table 8 and Figure 3. TempRetriever consistently outperforms baseline models and pretrained temporal models across validation and test sets. For example, VanillaDPR achieves a Top-1 accuracy of 43.10% and a Top-100 accuracy of 89.50% on the validation set, while DateAsToken slightly improves with a Top-100 accuracy of 89.00%. Among temporal models, BiTimeBERT performs the worst (Top-1: 27.70%, Top-100: 81.80%), highlighting its limitations in reasoning about implicit temporal information. TS-Retriever performs better, reaching a Top-1 accuracy of 39.60% and a Top-100 accuracy of 90.90%.
TempRetriever models, particularly TempRetriever FS, achieve state-of-the-art results, with a Top-1 accuracy of 44.80% and a Top-100 accuracy of 90.20%. TempRetriever VS achieves the highest Top-1 accuracy of 45.00%. Ranking quality metrics (Figure 3) further demonstrate TempRetriever ’s ability to prioritize relevant passages. For instance, at nDCG@1, VanillaDPR achieves 28.30%, which improves to 29.60% and 30.30% with TempRetriever FS and TempRetriever VS, respectively. This trend persists across all $k$-values, with TempRetriever FS achieving the highest scores, such as 45.09% at nDCG@100, compared to 43.78% for VanillaDPR and 29.29% for BiTimeBERT.
@@ 6.2. Addressing Different Question Types
We next explain how to construct a unified framework for handling different question types. Temporal retrieval systems must address diverse question types, including explicit temporal, implicit temporal, and non-temporal queries. As illustrated in Fig. 4, a query routing mechanism can be implemented to dynamically classify and assign questions to the appropriate retriever. Explicit temporal questions, which contain direct references to dates, are routed to TempRetriever, leveraging its ability to integrate temporal embeddings and fusion techniques.
Implicit temporal questions that lack explicit date mentions are routed to the query date prediction model, which predicts a temporal context based on the query’s content. The predicted date is then used to enhance the retrieval process through TempRetriever, aligning the inferred temporal context with the document timestamps. Finally, non-temporal queries, which do not involve temporal elements are directed to a general retriever like VanillaDPR. This simple routing framework optimizes retrieval performance across diverse query types by aligning each query with the most effective retriever.
@@ 7. RAG for Temporal Question Answering
The integration of TempRetriever into the Retrieval-Augmented Generation (RAG) system involves replacing the traditional retriever component with TempRetriever, which retrieves the top-k passages most relevant to a given query by considering both semantic and temporal relevance. These passages are then passed to a pre-trained large language model (LLM), such as Llama-3-8B (AI@Meta, 2024), for answer generation. The pipeline includes three main steps: encoding the query with its temporal context to create dense representations, ranking passages based on their combined semantic and temporal alignment, and generating answers using the retrieved passages as input for the LLM.
For the experiments, Llama-3-8B was used as the LLM, and its performance was evaluated on the ChroniclingAmericaQA and ArchivalQA datasets. The retriever retrieved top-1 passages for each query, which were fed into the LLM for supporting answer generation. Metrics such as Exact Match (EM), Recall (R), and Contains (Con) were used to compare the performance of RAG systems utilizing TempRetriever, VanillaDPR, and other baseline models.
Table 9 presents the results of the RAG framework for temporal question answering on explicit datasets. TempRetriever consistently outperformed other retrieval methods across all evaluated models and datasets. For example, when paired with Llama-3-8B, TempRetriever achieved an EM score of 16.90% and a recall of 28.52% on the validation set of ChroniclingAmericaQA. These scores were significantly higher than those obtained by the DateAsTag and DateAsToken models, which, while showing improvements over VanillaDPR, fell short of TempRetriever ’s performance. The DateAsToken model achieved an EM score of 16.57% on the same dataset, slightly below TempRetriever.
Across different LLMs, TempRetriever provided superior inputs for the generation task, as reflected in its higher metrics. For instance, with Llama-3-8B, TempRetriever achieved an EM score of 19.65% on ChroniclingAmericaQA and 45.21% on ArchivalQA, demonstrating its robustness in handling temporally-sensitive queries. These results underline TempRetriever’s ability to deliver temporally-aligned and contextually-relevant passages, enabling more accurate and effective answer generation in RAG systems.
@@ 8. Conclusion
In this paper, we introduced TempRetriever, a novel framework for Temporal Information Retrieval that integrates temporal information into dense passage retrieval using fusion techniques. By incorporating temporal signals directly into query and document representations, TempRetriever bridges the gap between semantic and temporal alignment, achieving superior performance across multiple retrieval metrics. Experiments on ArchivalQA and ChroniclingAmericaQA datasets demonstrate that TempRetriever significantly outperforms state-of-the-art baselines, including VanillaDPR, DateAsToken, and DateAsTag models, as well as temporal retrieval models like BiTimeBERT and TS-Retriever.
We also evaluated TempRetriever in a Retrieval-Augmented Generation (RAG) setup, where it provided high-quality inputs to large language models, leading to improved answer generation in temporal question answering. Looking forward, we plan to enhance TempRetriever’s handling of implicit temporal queries by integrating more advanced query date prediction models.